Given the proportion a/b=8/15, what ratio completes the equivalent proportion a/8=?

Answer by ad_alta(240)

1) a/b=8/15 [divide both sides by 8]
2) a/(8b)=1/15 [multiply both sides by b]
3) a/8=b/15